# Creative Problem Solving in Large Language and Vision Models – What Would it Take?

### v1

@@ III-A3 Transformational Creativity
These approaches involve transforming existing conceptual spaces to produce new ones. Transforming conceptual spaces can involve “altering existing rules[1], p.216. One way of transforming a model’s embedding space involves fine-tuning or training [40]. However, additional insight into transformational creative problem solving comes from prior work in CC, that describes creative problems as those with a poorly defined structure where a solution is not immediately apparent [41]. And in such cases, “… re-representation being the process which transforms an ill-structured problem into a well-structured one with direct inference to a problem solution[41], p.1. The notion of “re-representing” or “redefining” the problem can be best captured in the input prompts provided to an LLVM. This most closely connects to prompt engineering and in-context learning (ICL).
@@
Prompt engineering augments LLVMs with task specific hints, called prompts, to adapt the LLVM to new tasks [42]. Relatedly, in-context learning is a prompting method that provides the LLVM with instructions for solving a new task without requiring additional training. Prior work has shown that in-context learning and gradient-based optimization are equivalent [43], thus connecting ICL to training or fine-tuning.
@@
Extension of transformational creativity to LLVMs: Task re-representations for creative problem solving, through prompting or ICL, has not been well explored within ML. Prompt engineering and ICL is a challenging task, since model performance depends strongly on the chosen prompts [44], further compounded by the fact that creative problems are inherently poorly defined [41]. However, useful insights can be derived from CC literature. For instance, regarding problems that require creatively re-purposing objects, the Object-replacement-object-composition (OROC) framework [45] illustrates re-representations of tasks, that can be translated into prompts. The paper defines three types of creative tasks involving objects, and their task re-representations [45], p.16:
@@
For task re-representation, affordances can refer to object properties that are relevant to the task, e.g., in some cases the shape may be relevant and in other cases, the material [45]. Within LLVMs, the affordances $Af_{X}$ or $Af_{Y}$ can be defined via natural language, or other modalities such as images. In the following section, we present preliminary experiments on using LLVMs for object replacement, with prompts that are inspired by the above task re-representations. However, an in-depth application of these re-representations as defined in CC to the field of ICL in LLVMs remains an open question.
III-A4 Summary
In the previous sections, we drew parallels between Boden’s three forms of creativity and approaches in LLVMs, emphasizing how principles from CC can potentially help enable creative problem solving skills in these models.
Integration with task planning: Given the three methods, we see that transformational and combinational approaches may be especially aligned with LLVMs for high-level task planning. In contrast, exploratory methods may be better suited for low-level planning, e.g., trajectory generation.
Creative problem solving as a combination of the three methods: An effective approach to creative problem solving may require all the three methods described in this section. While papers have explored chaining of LLMs within frameworks (often via prompts) [46, 47], the individual LLMs themselves do not exhibit the characteristics described here. Existing frameworks in CC have shown that achieving creative problem solving would take a combination of all three methods, each of which is triggered in different contexts [41]. This presents potential opportunities for ML approaches that develop frameworks using multiple LLVMs. In particular, using CC frameworks such as “CreaCogs[45] as a start, can be highly beneficial for productive developments in ML.
Model Overall acc. % (no creativity)
@@ III-A4 Summary
TABLE I: Accuracy of the models in predicting the nominal use of objects with no creativity involved.
III-B What information should the LLVM embeddings be augmented with?
In the previous section, we discussed three methods for augmenting LLVM embedding spaces. In this section, we explore the question: “What information should be targeted by the three methods when augmenting the embedding space for creative problem solving?”. In the previous section, we discussed this in the context of OROC. According to the OROC framework [45], information about object affordances could enable models to re-represent the task, such that the solution becomes evident. We propose a small experiment to validate whether the principles of transformational creativity from OROC are useful to LLVMs. We note that creativity can occur in various contexts, e.g., creatively solving a math problem or creatively playing a chess move, each of which would require different information. However, to facilitate the discussion in this paper, we focus on solving tasks that require innovatively replacing missing objects (Task #1 of OROC).
III-B1 Experiment Setup
@@
For each group, we create test sets with 4 objects each, chosen from a set of RGB images of 16 objects (Appendix Figure 7). We create 10 such test sets per core object (total 50 samples per model). Each test set only includes one ground truth object, along with three other random objects that will not suit as an appropriate replacement. In the nominal group, the ground truth is the actual object itself. In the object replacement group, the replacements are chosen based on self-assessment of the authors as (core object $\xrightarrow{}$ replacement): “Scoop” $\xrightarrow{}$ “Bowl”; “Hammer” $\xrightarrow{}$ “Saucepan”; “Spatula” $\xrightarrow{}$ “Knife”; “Toothpick” $\xrightarrow{}$ “Safety pin”; “Pliers” $\xrightarrow{}$ “Scissors”. For each test case, we pass the images in the test set along with a prompt. We record whether the ground truth object image was chosen by the model for the prompt (i.e., assigned highest output probability) [^3] CLIP generates probabilities that given images correspond to a text. ViLT responds with a text, and we evaluate if the model responded “yes” with a high probability for the ground truth..
The nominal group is subjected to one type of prompt: “Can this object be used as a $\bigl{\langle}core\_object\bigl{\rangle}?$”. In the object replacement group, each test case is subjected to four prompts:
@@
    Baseline (regular) prompt: The same prompt as used in the nominal cases to obtain a baseline.
@@
    Prompt prepended with task and affordance information: the prompt includes additional information about the task and object affordance.
@@
Case #2 aligns with task re-representations of OROC, and we explore cases #3 and #4 for comparison. We formulate our affordance prompts as brief versions of OROC’s task re-representations. According to [45] affordances can be defined using shape features, which we apply to the prompts here. The full set of prompts is shown in Appendix Table II. The models that we explore include versions of CLIP [37], and ViLT [36] obtained from HuggingFace. We use different model sizes (Base, Large, Huge) and patch sizes (14, 16, 32). The open-source code for reproducing our experiment results (including our dataset and test cases) is available at: https://github.com/lnairGT/creative-problem-solving-LLMs.

### v2

@@ 4.1.3 Transformational Creativity
Transformational approaches involve transforming existing conceptual spaces to produce new ones. Transforming conceptual spaces can involve “altering existing rulesBoden (1998), p.216. One way of transforming a model’s embedding space involves fine-tuning or training Franceschelli and Musolesi (2023). However, additional insight into transformational creative problem solving comes from prior work in CC, that describes creative problems as those with a poorly defined structure where a solution is not immediately apparent Olteteanu (2014). And in such cases, “… re-representation being the process which transforms an ill-structured problem into a well-structured one with direct inference to a problem solutionOlteteanu (2014), p.1. The notion of “re-representing” or “redefining” the problem can be best captured in the input prompts provided to an LLVM. This most closely connects to prompt engineering and in-context learning (ICL).
@@
Prompt engineering augments LLVMs with task specific hints, called prompts, to adapt the LLVM to new tasks Gu et al. (2023). Relatedly, in-context learning is a prompting method that provides the LLVM with instructions for solving a new task without requiring additional training. Prior work has shown that in-context learning and gradient-based optimization are equivalent Von Oswald et al. (2023), thus connecting ICL to training or fine-tuning.
@@
Extension of transformational creativity to LLVMs: Task re-representations for creative problem solving, through prompting or ICL, has not been well explored within ML. Prompt engineering and ICL is a challenging task, since model performance depends strongly on the chosen prompts Rubin et al. (2021), further compounded by the fact that creative problems are inherently poorly defined Olteteanu (2014). However, useful insights can be derived from CC literature. For instance, regarding problems that require creatively re-purposing objects, the Object-replacement-object-composition (OROC) framework Olteţeanu and Falomir (2016) illustrates re-representations of tasks, that can be translated into prompts. The paper defines three different types of creative tasks involving objects, and their task re-representations as (from Olteţeanu and Falomir (2016), p.16):
@@
For task re-representation, affordances can refer to object properties that are relevant to the task, e.g., in some cases the shape may be relevant and in other cases, the material Olteţeanu and Falomir (2016). Within LLVMs, the affordances $Af_{X}$ or $Af_{Y}$ can be defined via natural language, or other modalities such as images. In the following section, we present preliminary experiments on using LLVMs for object replacement, with prompts that are inspired by the above task re-representations. However, an in-depth application of these re-representations as defined in CC to in-context learning in LLVMs remains unexplored.
4.1.4 Summary
In the previous sections, we drew parallels between Boden’s three forms of creativity and approaches in LLVMs, further emphasizing how principles from CC can potentially help enable creative problem solving skills in these models.
Integration with task planning: Given the three methods, we see that transformational and combinational approaches may be especially aligned with LLVMs for high-level task planning. In contrast, exploratory methods may be suited to low-level planning, e.g., trajectory generation.
Creative problem solving as a combination of the three methods: An effective approach to creative problem solving may require all the three methods described in this section. While papers have explored chaining of LLMs within frameworks (often via prompts) Karpas et al. (2022); Ling et al. (2023), the individual LLMs themselves do not exhibit the characteristics described here. Existing frameworks in CC have shown that achieving creative problem solving would take a combination of all three methods, each of which is triggered in different contexts Olteteanu (2014). This presents potential opportunities for ML approaches that develop frameworks using multiple LLVMs, e.g., extending CC frameworks such as “CreaCogsOlteţeanu and Falomir (2016) can be highly beneficial for productive developments in ML.
Model Acc. % (no creativity)
@@ 4.1.4 Summary
LLaVA 98.0%
Table 1: Accuracy of the models in predicting the nominal use of objects with no creativity involved.
4.2 What information should LLVM embeddings be augemented with?
In the previous section, we discussed three methods for augmenting LLVM embedding spaces. In this section, we explore the question: “What information should be targeted by the three methods when augmenting the embedding space for creative problem solving?”. In the previous section, we discussed this in the context of OROC. According to the OROC framework Olteţeanu and Falomir (2016), information about object affordances could enable models to re-represent the task, such that the solution becomes evident. We propose a small experiment to validate whether the principles of transformational creativity from OROC are useful to LLVMs. We note that creativity can occur in various contexts, e.g., creatively solving a math problem or creatively playing a chess move, each of which would require different information. However, to facilitate the discussion in this paper, we focus our scope on tasks that require innovatively replacing missing objects (OROC Task #1).
Note on embeddings vs. concepts: Our work connects “conceptual spaces” (or “concepts”) as defined in Computational Creativity literature, to “embedding spaces” (or “embeddings”) as defined in typical LM literature. We use “concepts” and “embeddings” interchangeably in this context. We make this connection to note that existing methods in Computational Creativity that operate on conceptual spaces translate to ML algorithms that operate on the LM’s embedding space. In this section, we connect the concept of “affordances” to the “embeddings” of the LLVMs in our experiments. Our goal is to show how the model can be prompted via an approach inspired by transformational creativity, to connect affordances of two seemingly distinct objects, e.g., a bowl and a spoon that appear distinct, but share the containability affordance.
4.2.1 Experiment Setup
@@
For each group, we create test sets with 4 objects each, chosen from a set of RGB images of 16 objects (Appendix Figure 3). We create 10 such test sets per core object (total 50 samples per model). Each test set only includes one ground truth object, along with three other random objects that will not suit as an appropriate replacement. In the nominal group, the ground truth is the actual object itself. In the object replacement group, the replacements are chosen based on self-assessment of the authors as (core object $\xrightarrow{}$ replacement): “Scoop” $\xrightarrow{}$ “Bowl”; “Hammer” $\xrightarrow{}$ “Saucepan”; “Spatula” $\xrightarrow{}$ “Knife”; “Toothpick” $\xrightarrow{}$ “Safety pin”; “Pliers” $\xrightarrow{}$ “Scissors”. For each test case, we pass the images in the test set along with a prompt. We record whether the ground truth object image was chosen by the model for the prompt (i.e., assigned highest output probability) [^2] CLIP generates probabilities that given images correspond to a text. ViLT and LLaVA respond with a text, and we evaluate if the model responded “yes” with a high probability for the ground truth..
The nominal group is subjected to one type of prompt: “Can this object be used as a $\bigl{\langle}core\_object\bigl{\rangle}?$”. In the object replacement group, each test case is subjected to four types of prompts:
@@
    Baseline (regular) prompt: Same prompt as used in the nominal cases to obtain a baseline.
@@
    Prompt prepended with task and affordance information: the prompt includes additional information on the task and object affordance.
@@
Case #2 aligns with task re-representations of OROC, and we explore cases #3 and #4 for comparison. We formulate our affordance prompts as brief versions of OROC’s task re-representations. According to Olteţeanu and Falomir (2016) affordances can be defined using shape features, which we apply to the prompts here. The full set of prompts is shown in Appendix Table 2. The models that we explore include versions of CLIP Radford et al. (2021), LLaVA Liu et al. (2024), and ViLT Kim et al. (2021) obtained from HuggingFace. We use different model sizes (Base, Large, Huge) and patch sizes (14, 16, 32). The open-source code for reproducing our experiment results (including our dataset and test cases) is available at: https://github.com/lnairGT/creative-problem-solving-LLMs. Appendix C includes more details on the experiments.
4.2.2 Results
In Table 1, we see the performances of the different models in the nominal test group, where the object requires no creative replacement. The models perform $>90\%$ in such cases (except for ViLT). In Figure 2, we see the performances (accuracy shown on a $0.0-1.0$ scale) of the models in the object replacement test cases, where the object requires a creative replacement. For reference, a model that randomly picks an object achieves about 30% overall accuracy. Figure 2 shows average accuracies for the different prompting strategies across random test sets. From Table 1 to Figure 2 (“regular”), the models perform poorly when they need to creatively reason about object replacements, highlighting their limitation. Comparing the “Regular” tab in Figure 2 to “Affordance”, we see a general improvement in model performances, when object affordance information is provided, consistent with description of the OROC framework Olteţeanu and Falomir (2016). However, information about the task (Figure 2, “Task” ) leads to mostly detrimental results. Information about task and affordances (Figure 2, “Task + Affordance”) does not lead to substantial improvements either, and is also detrimental in certain cases. We note that there is quite a variance in performances across the different models, which may be partially attributed to the original training datasets of the models. These observations warrant further exploration beyond the scope of this paper. Appendix D includes a detailed, class-wise breakdown of the results.
Figure 2: Object replacement group: Average accuracies and standard deviations of the models across ten different sets of randomly chosen objects.
4.2.3 Summary
While the experiments that we conducted are only preliminary, they offer some validity that the extension of principles in Computational Creativity can help overcome limitations of LLVMs in creative problem solving. The notion of task re-representation via improved prompting warrants further investigation in LLVMs, with regards to how the prompts can be generated automatically based on the creative task.
The models used in our experiments have all been trained jointly in visual and text domains. Multi-modal prompting capabilities may be useful for achieving creative problem solving. It can be quite challenging to describe affordances in words (example of “hammers” in our tests) and they may be better described through other means, e.g., images or depth maps or spectral data for material properties Erickson et al. (2020). This would require application of multi-modal LLVMs that can process a variety of data types Girdhar et al. (2023); Han et al. (2023). Computational creativity can offer insights into meaningful representations of these different modalities that would help achieve creative problem solving, e.g., whether object material or shape matters more for one task vs. another Olteţeanu and Falomir (2016).
5 Evaluation of Creativity
An important discussion in the context of creative problem solving is, how can creative problem solving be evaluated?. Prior work has proposed that creativity necessitates both novelty and value Boden (1998); Runco and Jaeger (2012), where the former guarantees that the generated outputs of a creative process are original, and the latter ensures that the generated outputs are useful. In the context of CPS, novelty refers to the discovery of new concepts (as defined in section 3.1), whereas value insists that the newly discovered concepts successfully solve the task. Hence, benchmarks for CPS should specifically evaluate how the task was solved (novelty and value) rather than the typical ML evaluation of whether the task was successful or not (value only). Some existing approaches that make this distinction describe problem settings that can be used to measure CPS skills of LLMs through the implicit integration of novelty and value measurements Tian et al. (2023); Naeini et al. (2023); Bisk et al. (2020); Talmor et al. (2022). In Tian et al. (2023), the authors create a dataset of 1600 real-world problems that necessarily involve creative reasoning abilities. Their proposed benchmark involves identifying novel approaches that can accomplish the given task (value). Similarly, in Naeini et al. (2023), the authors introduce the Only-Connect-Wall (OCW) dataset to measure CPS capabilities of LLMs. The authors in Bisk et al. (2020) explore physical commonsense reasoning that is more generally applicable, beyond object-based creative problems. The authors introduce Physical Interaction: Question Answering, or PIQA consisting of 16,000 QA pairs where each question is paired with two possible common-sense solutions with a ground truth. In Talmor et al. (2022), the authors introduce CommonSenseQA 2.0 (CSQA2) dataset consisting of both object-based and non-object based creative problems. The dataset consists of 14,343 questions distributed across 1,868 distinct topics. Currently, to the best of our knowledge, there are no standard benchmarks available to measure CPS skills of VLMs, although our preliminary experiments show one way to measure this using the task of object substitution.
6 Conclusion and Future Work
In this paper, we argued that an effective approach for enabling creative problem solving – currently a key limitation of LLVMs – should derive from Computational Creativity literature. To emphasize this at each juncture, we discussed the specific principles from CC that can be extended to achieve creative problem solving in LLVMs, describing the potential for further research with these insights. It is rare to see special tracks or workshops targeted at Computational Creativity within more prestigious ML conferences. These programs typically focus on creative artifact generation and art (such as the NeurIPS Workshop on Machine Learning for Creativity and Design NeurIPS (2022) or the recent tutorial at EMNLP on Creative Natural Language Generation Chakrabarty et al. (2023)), but do not discuss CPS, thus failing to bridge the gap between CC and ML. We hope to see a deeper integration of the CC communities at such strong ML venues. We hope to encourage the reader to view creative problem solving and ML holistically, through the lens of Computational Creativity.
7 Limitations
Literature outside of Computational Creativity that enables CPS is unexplored: Our paper predominantly focuses on CC literature. This work does not cover literature beyond CC that can potentially inform creative problem solving in LLVMs. Although CC literature broadly encompasses psychology, neuroscience and philosophy, our future work seeks to explore specific literature within these sub-domains and discuss their applicability to creative problem solving and ML.
Lack of an explicit creative problem solving algorithm for LLVMs: Since the scope of our work aligns with a position paper, we have not focused on developing a concrete algorithm for creative problem solving in LLVMs. The prompting strategies explored in our preliminary experiments are manually specified, and our work does not elaborate on how these prompts may be automatically discovered. While our paper seeks to address some of the key gaps that prevent the application of CC literature to ML, there are still several unanswered questions when it comes to the practical implementation of an ML approach: e.g., what is a good representation for concepts that facilitate creative problem solving (symbolic, non-symbolic, or hybrid)? What is a good problem formulation for a given creative problem solving task (planning or learning)? etc. However, these questions are not directly answered within the scope of our work.
8 Ethical Considerations
The authors do not have specific ethical considerations to be highlighted with respect to this work.
Appendix A Alternate Definitions of Creative Problem Solving
Prior work by Olteţeanu Olteteanu (2014) defines CPS from an object affordance perspective, where affordances broadly refer to action possibilities for objects, e.g., cups are pour-able and doors are open-able. The authors in Olteteanu (2014) define creative problems as nominal problem solving tasks that have a poor representational structure, and as “the ability of a cognitive, natural, or artificial system to use new objects to solve a problem, other than the ones that have been stored in its memory as tools for that specific purpose (if any), or to create those objects by putting together objects or parts of objects the system has access to. Depending on the problem, objects can be either physical or abstract/informational (concepts, problem templates, heuristics or other forms of representations)”. However, this definition is primarily object-creativity centered, and does not cover a wider range of creative problems.
Follow-up work by Sarathy and Scheutz Sarathy and Scheutz (2018), define “Macgyver-esque” creativity as a planning task that involves “generating, executing, and learning strategies for identifying and solving seemingly unsolvable real-world problems”. They introduce the “MacGyver Problem” (MGP) as a planning problem with an unreachable goal state. Through the modification of the agent’s domain knowledge (through domain expansion and domain contraction), the agent must discover new information and incorporate it into its existing domain knowledge, allowing the agent to accomplish the task. The domain expansion and contraction processes align with the divergent-convergent model of creative problem solving Guilford (1967). The definition of an MGP aligns well with the formulation of planning problems in ML, but less with learning or hybrid planning-learning approaches.
Appendix B Alternate theories on creative problem solving and their applications to ML
While there is exhaustive literature regarding theories on general creativity, we focus specifically on creative problem solving, with three well received works: Divergent-Convergent Thinking Guilford (1967), Explicit-Implicit Interaction Theory Hélie and Sun (2010), and the Creative Systems Framework Wiggins (2006). We discuss their applicability to ML in addition to the literature discussed in the main body of this paper. Our goal in this section is to further widen the discussion on integrating CC and ML to achieve creative problem solving in LLVMs, with additional literature.
B.0.1 Divergent-Convergent Thinking
In Guilford (1967), the authors discuss the notion of “divergent-convergent” thinking. Divergent thinking or “divergent-production” (DP) abilities involve a more open-ended generation of a variety of ideas, whereas convergent thinking focuses on applying specific ideas to solve the problem.
Applicability to CPS in LLVMs: Prior work by Tian et al. (2023) have demonstrated the applicability of “divergent-convergent” thinking towards solving Macgyver problems. Similar in spirit to our experiments with VLMs in Section 4.2.1, the authors prompt LLMs with descriptions of objects to enable the LLMs to reason about solving the task. Their work is, to the best of our knowledge, the only direct example demonstrating the value of CC literature in enabling CPS in LLMs.
B.0.2 Explicit-Implicit Interaction Theory
In Hélie and Sun (2010), the authors introduce the Explicit-Implicit Interaction (EII) theory, building upon the seminal work in Wallas (1926), that describes four stages of creativity: Preparation, incubation, illumination (i.e., insight), and verification. Preparation refers to the initial stage of searching in many different directions, which may fail to find a solution (i.e., impasse) in case of ill-defined problems (as is the case with CPS). Following an impasse, the incubation phase begins, where attention is not devoted to solving the problem. Over a period of time, illumination is the manifestation of the solution to the problem within the conscious thought (i.e., “Aha” moment). Finally, verification involves using deliberative thinking to assess if the solution indeed solves the problem.
Applicability to CPS in LLVMs: The authors in Hélie and Sun (2010) incorporate the four stages via a concrete computational method into the CLARION cognitive architecture. Prior work has also introduced a CPS framework for ML approaches inspired by the four stages Gizzi et al. (2022). In their work, “preparation” aligns with problem formulation, either task learning or planning. Incubation and illumination aligns with knowledge representation (symbolic, non-symbolic, or hybrid), and knowledge manipulation (functions that manipulate the conceptual space). Lastly, verification aligns with evaluation (via simulation, real-world platforms, or benchmarks). Although these works do not explicitly cover LLVMs and related algorithms, they demonstrate the value of integrating CC literature in ML, and can serve as useful starting points for ML approaches towards creative problem solving in LLVMs.
B.0.3 Creative Systems Framework
In Wiggins (2006), the author expands on Boden’s levels further in the context of a framework that formalizes creative systems. The paper defines: a) creative system, b) creative behavior, c) novelty, and d) value. The paper also discusses formalized notion of a universe of possibilities, and conceptual spaces. Crucially, the work describes the characteristics of a creative agent, that can help distinguish modes of failures within a creative system, namely: a) hopeless uninspiration – where there are no valued concepts within the universe; b) conceptual uninspiration – where there are no valued concepts within the conceptual space of the agent; and c) generative uninspiration – where an agent is unable to find a valued concept owing to the specific method (e.g., search) employed.
Applicability to CPS in LLVMs: While the discussion of novelty, value and conceptual spaces in Wiggins (2006) aligns with our descriptions in Section 4, the different modes of uninspiration offers potential ways to assess failure modes in LLVMs. This allows agents to distinguish between systems where creative problem solving is not possible (hopeless uninspiration), as compared to systems where the conceptual space or the methodology for searching the conceptual space, may be at fault (conceptual or generative uninspiration). Although this approach has not been expanded in existing literature, it presents a promising direction for an evaluation framework that can distinguish CPS from non-CPS problems.
B.1 A potential link between creative problem solving and general intelligence
Existing literature hints at a potential link between creative problem solving and Artificial General Intelligence (AGI) - systems that are broadly capable of solving almost all tasks that humans can Shevlin et al. (2019). For instance, in Moruzzi (2020), p.85., the author argues that there exists a strong correlation between creativity and AGI: “… features that systems need to develop in order to achieve general intelligence are aspects that they need to possess also to earn the attribute creative”. In Goertzel (2014), the author compiles a list of competencies deemed essential for achieving AGI, including creative capacities like “conceptual invention” and “creative constructive play with objects”. The processes of “insight” or “incubation” often associated with creative problem solving Hélie and Sun (2010); Gilhooly (2016) is also considered important for AGI Ventura (2014). Taken together, it is likely that any promising vision of AGI would be incomplete without creative problem solving.
Alongside the heavy ongoing discussion of AGI surrounding LLVMs Bubeck et al. (2023); Fei et al. (2022); Ma et al. (2023); Xi et al. (2023); Moor et al. (2023); Grudin and Jacques (2019), there is often little to no discussion of creative problem solving or Computational Creativity within mainstream ML. As described in Moruzzi (2020), p.96, “The investigation on the nature of creativity and on how it manifests itself not only in human but also in animal and artificial systems should, thus, not be intended as a niche discussion but, rather, as a fundamental research which can lay the foundations for further studies in artificial intelligence and its relation to humans”. We hope that this work will encourage discussions of creative problem solving and Computational Creativity alongside discussions on AGI.
Appendix C Experiment Settings
Prompt type Prompt
Regular
“can this object be used as a scoop?”
“can this object be used as a hammer?”
“can this object be used as a spatula?”
“can this object be used as a toothpick?”
“can this object be used as pliers?”
Affordance
“scoops must be concave and hollow. can this object be used as a scoop?”
“hammers must be heavy and have a handle attached to a cylinder at the end.
can this object be used as a hammer?”
“spatulas must have a handle attached to a flat surface at the end.
can this object be used as a spatula?”
“toothpicks must have a pointed tip. can this object be used as a toothpick?”
“pliers must have two-prongs. can this object be used as pliers?”
Task
“scoops can transfer beans from one jar to another jar. can this object be
used as a scoop?”
“hammers can hit a nail into the wall. can this object be used as a hammer?”
“spatulas can spread butter onto a pan. can this object be used as a spatula?”
“toothpicks can pick food caught between the teeth. can this object be used
as a toothpick?”
“pliers can grab a coin. can this object be used as pliers?”
Task and affordance
“scoops can transfer beans from one jar to another jar. scoops are concave
and hollow. can this object be used as a scoop?”
“hammers can hit a nail into the wall. hammers have a handle attached to a
cylinder at the end. can this object be used as a hammer?”
“spatulas can spread butter onto a pan. spatulas have a handle attached to a
flat surface at the end. can this object be used as a spatula?”
“toothpicks can pick food caught between the teeth. toothpicks have a
pointed tip. can this object be used as a toothpick?”
“pliers can grab a coin. pliers have two-prongs. can this object be used as
pliers?”
Table 2: Prompts (across 4 groups) used in the experiment
Figure 3: Complete test set of objects used in the experiments.
C.1 Data: Test images
Figure 3 shows the test set of 16 RGB images of objects used for the object substitution task. From the shown image dataset, we create test sets with 4 objects each, chosen from the set of 16 object images. We create 10 such test sets per core object (total 50 samples per model). Each test set only includes one ground truth object, along with three other random objects that will not suit as an appropriate replacement. In the nominal group, the ground truth is the actual object itself. In the object replacement group, the replacements are chosen based on self-assessment of the authors as (core object $\xrightarrow{}$ replacement): “Scoop” $\xrightarrow{}$ “Bowl”; “Hammer” $\xrightarrow{}$ “Saucepan”; “Spatula” $\xrightarrow{}$ “Knife”; “Toothpick” $\xrightarrow{}$ “Safety pin”; “Pliers” $\xrightarrow{}$ “Scissors”.
C.2 Model: Checkpoints
For all the models, we use pre-trained HuggingFace checkpoints, with no additional training or fine-tuning. The models are of different architecture sizes and patch sizes: “CLIP-B-32” uses the “openai/clip-vit-base-patch32” which is a base model with a patch size of 32. “CLIP-B-16” uses “openai/clip-vit-base-patch16” – a base model with patch size of 16. “CLIP-L-14” uses “openai/clip-vit-large-patch14” – a large model with patch size of 14. “CLIP-H-14” uses “laion/CLIP-ViT-H-14-laion2B-s32B-b79K” which is a “huge” model, with a patch size of 14. This model is trained with the 2 billion sample English subset of LAION-5B. For LLaVA, we use the “llava-hf/llava-1.5-7b-hf” with 7B parameters, version 1.5. Lastly, “VILT-B-32” uses “dandelin/vilt-b32-finetuned-vqa” trained for visual question answering. However, there is limited data available on HuggingFace regarding the model.
C.3 Prompts used in testing
In this section, we discuss the prompts used in the different testing conditions (see Table 2). We explore four classes of prompts for the creative object substitution task: “Regular”, “Affordance”, “Task” and “Task and affordance”. Regular prompts involve a direct prompt as to whether a given object will suffice as a substitute for the missing object. Affordance prompts, adds information about the desired affordances that are essential for replacing the missing object. Task prompts adds additional information on the task to be performed as context for whether a given object can be used as replacement for the missing object. Lastly, task and affordance prompts combine the task and object affordance information within the prompt.
C.4 Testing Procedure
For each test case, we pass the images in the test set along with a prompt belonging to one of the four classes described in Table 2. We record whether the ground truth object image was chosen by the model for the prompt (i.e., assigned highest output probability). CLIP generates probabilities that given images correspond to a text. ViLT responds with a text, and we evaluate if the model responded “yes” with a high probability for the ground truth.
C.5 Testing Infrastructure
We used NVIDIA-A100 GPUs to run the evaluation. However, the models are not too large and we have tested and confirmed that the code can be executed on CPU only as well.
Appendix D Continued Experiment Results
In this section, we show the class-wise breakdown of the different models for the different prompting strategies (Figures 4 - 7). We note that “hammers” present a particularly challenging case for all the models, perhaps due to the fact that correlating affordance of a hammer to a saucepan textually is difficult. In contrast, all models with the augmented prompts typically perform well in the case of creatively replacing “toothpick” with “safety pin” – presumably indicating that specifying the relevant affordance textually in this case provides sufficient information. We repeated each experiment across multiple random seeds and found similar performances, showing that our general findings hold across different random cases. Generally, specifying object affordance information in the prompts leads to improved model performance.